OpenStudy (anonymous):

Factor completely: 2x2 + 2x - 24

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

na

OpenStudy (skullpatrol):

$$\Huge 2x^2 + 2x - 24$$

OpenStudy (anonymous):

i need it factored .

OpenStudy (skullpatrol):

Start by looking for the greatest common factor first.

OpenStudy (skullpatrol):

$$\Huge 2x^2 + 2x - 24$$ $$\Huge= 2x^2 + 2x - 2*12$$

OpenStudy (skullpatrol):

See anything common?

OpenStudy (anonymous):

2

OpenStudy (skullpatrol):

Factor it out please :)

OpenStudy (skullpatrol):

$$\Huge ab+ac=a(b+c)$$

OpenStudy (skullpatrol):

$$\Huge 2x^2 + 2x - 2*12=?$$